### v1

# Generative modelling with jump-diffusions

Abstract
Score-based diffusion models generate samples from an unknown target distribution using a time-reversed diffusion process. While such models represent state-of-the-art approaches in industrial applications such as artificial image generation, it has recently been noted that their performance can be further improved by considering injection noise with heavy tailed characteristics. Here, I present a generalization of generative diffusion processes to a wide class of non-Gaussian noise processes. I consider forward processes driven by standard Gaussian noise with super-imposed Poisson jumps representing a finite activity Lévy process. The generative process is shown to be governed by a generalized score function that depends on the jump amplitude distribution. Both probability flow ODE and SDE formulations are derived using basic technical effort, and are implemented for jump amplitudes drawn from a multivariate Laplace distribution. Remarkably, for the problem of capturing a heavy-tailed target distribution, the jump-diffusion Laplace model outperforms models driven by $\alpha$-stable noise despite not containing any heavy-tailed characteristics. The framework can be readily applied to other jump statistics that could further improve on the performance of standard diffusion models.
1 Introduction
Generative diffusion models transform a target distribution into a simple distribution by successively adding noise and then reversing this process. By sampling from the simple distribution, the reversed (or denoising) process is thus able to generate a new target sample representing, e.g., a new image of a tree. Starting from this seminal concept [15], two main implementations have been developed, where, roughly speaking, the denoising process is either formulated in discrete (time) steps and represented by a Markov chain, for example implemented in denoising score-matching (DSM) [19] and denoising diffusion probabilistic models (DDPMs) [15, 6], or implemented in the form of a stochastic differential equation (SDE) in continuous time [18]. In both cases, the underlying noise process is typically taken to have Gaussian characteristics, which facilitates both theoretical treatments and computational implementations.
However, from a performance perspective, it is not a priori obvious that Gaussian noise represents the optimal choice in specific applications. Indeed, the question of how the performance of generative diffusion models can be improved by changing the characteristics of the underlying noise has recently attracted interest in the research community. In [11] it has been shown that using noise from a Gamma distribution in DDPM improves results for image and speech generation. In the SDE approach of [12], Gaussian noise has been replaced by noise from fractional Brownian motion, which achieves greater pixel-wise diversity and enhanced image quality. The effect of $\alpha$-stable noise has been investigated both in the SDE formulation [20] and DDPM [14], and has been shown to allow for faster and more diverse sampling. In [13] DDPM has been implemented with noise from a heavy-tailed Student-t distribution and shown to outperform standard diffusion models in heavy-tail estimation on high-resolution weather datasets. In [7] a Markovian jump model based on generator matching has been proposed, which achieves state-of-the-art performance in multi-modal protein generation.
In this work, I present a non-Gaussian generalization of the conventional SDE framework of diffusion models [18] that includes discontinuous jumps in the forward noising process. A similar idea has previously been explored in [20], but this approach is technically complicated and only tailored for a specific type of heavy-tailed noise. On the other hand, focusing on the mathematically much simpler class of jump-diffusion processes with finite activity, my framework represents a straightforward generalization of the Gaussian case and covers a wide-range of jump amplitudes.
The contributions of the present work are summarized as follows:
  • A score-based generative model for jump-diffusions as underlying noise process is presented, generalizing the work of [18]
  • Representations of the generative process in terms of a probability flow ODE and an SDE with jump-diffusion noise are derived. These are driven by a generalized score function $\neq\nabla\log\,p(\mathbf{x},t)$.
  • The detailed implementation for jump amplitudes drawn from a multivariate Laplace distribution is discussed. The model is shown to outperform methods based on heavy-tailed noise [14, 20] in capturing the tails of a heavy-tailed target distribution.
1.1 Relationship with other works
The present work is closely related to the work of [20], who first considered the case of jumps in the forward process in the Lévy-Ito model (LIM). While LIM only focuses on the special case of infinite activity $\alpha$-stable noise with $21<\alpha<$ 2, I consider finite activity Lévy noise given as a superposition of Gaussian noise with Poisson jumps drawn from an arbitrary normalizable distribution. The mathematical discussion is then much simplified and provides a pedagogical access to non-Gaussian generalizations of the conventional SDE framework. More details on the comparison with LIM are provided in Sec. C.2.
DDPMs with specific non-Gaussian noise distributions have been investigated in [11] (gamma), [14] ($\alpha$-stable), [13] (Student-t). DSM has been extended to generalized Gaussian noise with heavy tails in [5]. The work of [10] considers random walks with non-Gaussian increments that converge to standard diffusion models. All these Markov chain approaches are formally different even in the limit of continuous time, since the non-Gaussian increments act at every time step of the underlying process, which is different from the jump structure of a Lévy process whose jump times are Poisson distributed and thus intermittent.
In [7] a specific Markovian jump model has been proposed based on matching with a Gaussian probability path, i.e., the jump model generates the same conditional probability distributions as a Gaussian model, but has different sample paths. On the other hand, the finite activity Lévy model considered here is always characterized by non-Gaussian propagators.
In [2] a generative diffusion model is proposed that includes Poisson jumps between different dimensional spaces. The jumps only act in terms of creating/destroying dimensions and do not denoise the samples within a given dimension.
2 Generative score-based diffusion models
Generative diffusion models sample from an (unknown) target distribution $p_{\rm data}$ by modelling an inverse time process. Starting point in the SDE formulation is the forward diffusion process
$\dot{\mathbf{Y}}(t)=\mathbf{f}(\mathbf{Y},t)+g(t)\boldsymbol{\xi}(t)$ (1)
for an initial data distribution $\mathbf{Y}(0)\sim p_{\rm data}(\mathbf{x})$ with sample space $\mathbf{x}\in\mathbb{R}^{d}$. The noise is typically taken as Gaussian white noise $\boldsymbol{\xi}(t)=\boldsymbol{\xi}^{\rm G}(t)$ with covariance $<\xi^{\rm G}_{i}(t)\xi^{\rm G}_{j}(t^{\prime})>=D\delta_{ij}\delta(t-t^{\prime})$. The forward process is distributed as $\mathbf{Y}(t)\sim p(\mathbf{x},t)$, with the probability density function (PDF) or propagator $p(\mathbf{x},t)=\int\mathrm{d}\mathbf{x}^{\prime}p(\mathbf{x},t|\mathbf{x}^{\prime})p_{\rm data}(\mathbf{x})$, where $p(\mathbf{x},t|\mathbf{x}^{\prime})$ denotes the conditional PDF of the forward process for the initial condition $\mathbf{Y}(0)=\mathbf{x}^{\prime}$.
The generative process then starts from $\mathbf{X}(0)\sim p(\mathbf{x},T)$ and generates samples of the unknown $p_{\rm data}$ using time-reversal. The time-reversed process is given by [1]
$\dot{\mathbf{X}}(t)=-\mathbf{f}(\mathbf{X},T-t)+Dg^{2}(T-t)\nabla\log\,p(\mathbf{X},T-t)+g(T-t)\boldsymbol{\xi}^{\rm G}(t)$ (2)
where the central ingredient is the score function $\nabla\log\,p(\mathbf{x},t)$, which directs the $\mathbf{X}$-process towards $p_{\rm data}$, and contains the PDF $p$ of the forward process. Describing the backwards evolution in terms of a PDF $\mathbf{X}(t)\sim\tilde{p}(\mathbf{x},t)$ and starting from $\tilde{p}(\mathbf{x},0)=p(\mathbf{x},T)$, we have $\tilde{p}(\mathbf{x},t)=p(\mathbf{x},T-t)$, i.e., the score function ensures that $\mathbf{X}$ at any time $t$ is equal in distribution to $\mathbf{Y}$ at time $T-t$. The challenge of sampling the starting point $\mathbf{X}(0)\sim p(\mathbf{x},T)$ is simplified by choosing $\mathbf{f},g$ such that $p(\mathbf{x},t|\mathbf{x}^{\prime})$ converges for $t\to\infty$ to a well-defined stationary distribution $p_{\rm st}(\mathbf{x})$, for example a simple Gaussian. For sufficiently large $T$, we have then $p(\mathbf{x},T)\approx p_{\rm st}(\mathbf{x})$ from which $\mathbf{X}(0)$ is sampled.
The generative process can be equivalently described by the deterministic probability flow ODE [18]
$\dot{\mathbf{X}}(t)=-\mathbf{f}(\mathbf{X},T-t)+\frac{D}{2}g^{2}(T-t)\nabla\log\,p(\mathbf{X},T-t),$ (3)
which generates the same distribution $\tilde{p}(\mathbf{x},t)$ as Eq. (2) provided the same initial distribution $\mathbf{X}(0)\sim p(\mathbf{x},T)$ is used.
Of course, if $p_{\rm data}$ is unknown, the exact analytical form of $p(\mathbf{x},t)$ and thus the score function is also not known. However, the generative power of this approach stems from the fact that the score function can be estimated from the available data (samples from $p_{\rm data}$) and thus yields an approximation of the exact sampling processes Eqs. (2,3). One can show that the parameters $\boldsymbol{\theta}$ of a given parametric form $\mathbf{s}_{\boldsymbol{\theta}}$ for the score function parametrized for example by a neural network, can be estimated by optimizing the loss function
$J_{\rm G}(\boldsymbol{\theta})$ $=$ $\int_{0}^{T}\mathrm{d}t\,w(t)\mathbb{E}_{\mathbf{Y}(0)}\mathbb{E}_{\mathbf{Y}(t)|\mathbf{Y}(0)}\left[\|\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{Y},t)-\nabla\log\,p(\mathbf{Y},t|\mathbf{Y}_{0})\|^{2}\right]$ (4)
with weight function $w(t)$. Eq. (4) is called the denoising score matching loss function and highlights that the estimation of $\mathbf{s}_{\boldsymbol{\theta}}$ does not require knowledge of the exact $p_{\rm data}$, but can be performed from finite samples of $p_{\rm data}$ using Monte-Carlo integration.
As discussed in [18], this SDE formulation unifies widely-used generative modelling approaches based on Markov chains, which correspond to discrete time approximations of Eqs. (1,2). For example, discretizing $t,\mathbf{X}(t),g(T-t)$ over $[0,T]$ with time step $\Delta t$ and setting $\mathbf{f}=\mathbf{0}$, $D=1$ yields Eq. (2) as the iterative equation
$\mathbf{X}_{i+1}=\mathbf{X}_{i}+g_{i}^{2}\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{X}_{i},T-t_{i})\Delta t+g_{i}\sqrt{\Delta t}\,\mathbf{Z}_{i},$ (5)
where $\mathbf{Z}_{i}\sim\mathcal{N}(0,\mathbf{I})$ and $\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{X}_{i},T-t_{i})$ approximates $\nabla\log\,p(\mathbf{X}_{i},T-t_{i})$. Eq. (5) corresponds to SDM Langevin dynamics [16] that generates samples of $p_{\rm data}$ by gradient descent plus noise over different noise scales $g_{i}$. Likewise, one can recover DDPM by setting $\mathbf{f}(\mathbf{y},t)=-\frac{1}{2}\beta(t)\mathbf{y}$, $g(t)=\sqrt{\beta(t)}$ and discretizing Eq. (2) [18].
3 From diffusions to jump-diffusions
The aim of this work is to consider the case when the noise $\boldsymbol{\xi}$ in Eq. (1) is not only given as Gaussian white noise, but also contains non-Gaussian contributions in the form of additional discontinuous jumps that occur with a fixed rate and whose amplitudes are drawn independently from a given distribution. The general class of such processes with stationary and independent increments is known as Lévy processes and have a well-developed mathematical theory [4]. In the following, I focus on the class of finite activity Lévy processes, for which the jump amplitude distribution is normalizable, excluding the case of power-law or gamma distributed jump amplitudes. The mathematical theory is considerably simpler in this case and leads to a straightforward generalization of the conventional Gaussian case discussed in Sec. 2. The special case of pure power-law jumps corresponding to $\alpha$-stable noise with $21<\alpha<$ 2 has been discussed in detail in the LIM framework [20] using rigorous methods, see also Sec. C.2.
Including jumps, the noise in Eq. (1) is now expressed as
$\boldsymbol{\xi}(t)=\boldsymbol{\xi}^{\rm G}(t)+\sum_{j=1}^{N_{T}}\mathbf{A}_{j}\delta(t-\tau_{j}),$ (6)
where $\boldsymbol{\xi}^{\rm G}(t)$ is the Gaussian white noise as before and the jumps are described as follows: the number of jumps in the time interval $[0,T]$ follows a Poisson process with rate $\lambda$, i.e., $P(N_{T}=n)=\frac{(\lambda T)^{n}}{n!}e^{-\lambda T}$, the jump instances $\tau_{j}$ are uniform in $[0,T]$, and the jump amplitudes $A_{j}$ are drawn i.i.d. from a distribution $\rho$. Considering a finite time step $\Delta t$, one can show that the statistics of the noise increments $\Delta\boldsymbol{\xi}=\int_{t}^{t+\Delta t}\boldsymbol{\xi}(t^{\prime})\mathrm{d}t^{\prime}$ is then described by the characteristic function [4]:
$G_{\Delta\boldsymbol{\xi}}(\mathbf{k})=\left<e^{i\mathbf{k}{\cdot}\Delta\boldsymbol{\xi}}\right>=e^{-\psi(k)\Delta t},\qquad\qquad\psi(k)=\frac{D}{2}k^{2}-\lambda\phi(k)$ (7)
with $k=\|\mathbf{k}\|$ and
$\phi(k)=\int\mathrm{d}\mathbf{z}\left(e^{i\mathbf{k}{\cdot}\mathbf{z}}-1\right)\rho(z).$ (8)
contains the characteristic function of the jumps. A crucial constraint is that I assume isotropic jump distributions such that $\rho$ is a function of $z=\|\mathbf{z}\|$ only and, as a consequence, also $\phi$ is a function of $k=\|\mathbf{k}\|$ only.
I then obtain the following jump-diffusion generalizations of Eqs. (2,3,4), when the noise of Eq. (1) is specified by Eqs. (7,8):
  1. 1.
    Generalized score function: Instead of $\nabla\log\,p(\mathbf{x},t)$, the score function that governs the generative process is now given by $\frac{\nabla\mathcal{V}(\mathbf{x},t)}{p(\mathbf{x},t)}$, where $p(\mathbf{x},t)$ is the PDF of the forward process $\mathbf{Y}(t)$ with Fourier transform (FT) $\hat{p}(\mathbf{k},t)$ and $\mathcal{V}$ is expressed in terms of the inverse FT
    $\mathcal{V}(\mathbf{x},t)$ $=$ $\mathcal{F}^{-1}\left\{\hat{p}(\mathbf{k},t)\frac{\psi(k\,g(t))}{k^{2}}\right\}$ (9)
    Eq. (9) recovers the conventional score function for $\lambda=0$. Likewise, setting $\phi(k)=-\frac{a^{2}}{2}k^{2}$ in Eq. (7), i.e., additional Gaussian noise with intensity $D_{2}=\lambda a^{2}$ just recovers the conventional case with noise intensity $D\to D+D_{2}$.
  2. 2.
    Generative process: The jump-diffusion equivalent of the probability flow equation Eq. (3) is given by
    $\dot{\mathbf{X}}(t)=-\mathbf{f}(\mathbf{X},T-t)+\frac{\nabla\mathcal{V}(\mathbf{X},T-t)}{p(\mathbf{X},T-t)}$ (10)
    The PDF of the generative process can likewise be generated by an SDE generalizing Eq. (2)
    $\dot{\mathbf{X}}(t)=-\mathbf{f}(\mathbf{X},T-t)+2\frac{\nabla\mathcal{V}(\mathbf{X},T-t)}{p(\mathbf{X},T-t)}+g(T-t)\boldsymbol{\xi}(t)$ (11)
    For the derivation of Eqs. (9,10,11) see Appendix A
  3. 3.
    Denoising score matching loss function: The jump-diffusion equivalent of Eq. (4) is given by
    $J_{\rm JD}(\boldsymbol{\theta})$ $=$ $\int_{0}^{T}\mathrm{d}t\,w(t)\mathbb{E}_{\mathbf{Y}(0)}\mathbb{E}_{\mathbf{Y}(t)|\mathbf{Y}(0)}\left[\left\|\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{Y},t)-\frac{\nabla\mathcal{V}(\mathbf{Y},t|\mathbf{Y}_{0})}{p(\mathbf{Y},t|\mathbf{Y}_{0})}\right\|^{2}\right]$ (12)
    where
    $\mathcal{V}(\mathbf{x},t|\mathbf{x}^{\prime})$ $=$ $\mathcal{F}^{-1}\left\{\hat{p}(\mathbf{k},t|\mathbf{x}^{\prime})\frac{\psi(k\,g(t))}{k^{2}}\right\}$ (13)
    and $p(\mathbf{x},t|\mathbf{x}^{\prime})$ denotes the propagator of $\mathbf{Y}(t)$, Eq. (1), with FT $\hat{p}(\mathbf{k},t|\mathbf{x}^{\prime})$. Eq. (12) can be derived from the explicit score matching loss function following standard steps, see Appendix B
  4. 4.
    Discrete approximations: Eq. (11) can be straightforwardly discretized leading to jump-diffusion generalizations of Markov chain diffusion models. For example, the jump-diffusion equivalent of Eq. (5) with $\mathbf{f}=\mathbf{0}$ is given by
    $\mathbf{X}_{i+1}=\mathbf{X}_{i}+g_{i}^{2}\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{X}_{i},T-t_{i})\Delta t+g_{i}\sqrt{\Delta t}\,\mathbf{Z}_{i}+g_{i}\Delta\mathbf{J}_{i}$ (14)
    where $\mathbf{s}_{\boldsymbol{\theta}}$ is estimated by Eq. (12) and $\mathbf{Z}_{i}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. The jump increment $\Delta\mathbf{J}_{i}$ is generated by drawing $M\sim\text{Poisson}(\lambda\Delta t)$ and setting $\Delta\mathbf{J}_{i}=\sum_{j=1}^{M}\mathbf{A}_{j}$, where $\mathbf{A}_{j}\sim\rho$ i.i.d. One obtains likewise a jump-diffusion variant of DDPM by setting $\mathbf{f}(\mathbf{x},t)=-\frac{1}{2}\beta(t)\mathbf{x}$, $g(t)=\sqrt{\beta(t)}$.
One of the key insights here is that changing the noise type in the forward process leads to a modification of the score function for consistent denoising according to the correct time-reversal. In particular the DSM formulation Eq. (14) is conceptually different from an heuristic approach that simply samples Langevin dynamics with the standard score function $\nabla\log\,p(\mathbf{x},t)$ plus non-Gaussian noise. I now discuss the detailed implementation of the framework for a specific choice of $\mathbf{f},g$, and jump amplitude distribution.
3.1 The jump-diffusion Laplace model
In the Gaussian case, it is desirable to choose $\mathbf{f},g$ such that: (i) a closed analytical form of the propagator $p(\mathbf{x},t|\mathbf{x}^{\prime})$ is available, which facilitates the score function approximation via Eq. (4); and (ii) to be able to generate samples from $p_{\rm st}(\mathbf{x})$ efficiently. A common choice is thus an Ornstein-Uhlenbeck process as forward process, where $p(\mathbf{x},t|\mathbf{x}^{\prime})$ is Gaussian. In the jump-diffusion case, we need to consider the loss function Eq. (12) and thus also need to evaluate the term $\nabla\mathcal{V}(\mathbf{x},t|\mathbf{x}^{\prime})$. As we will see, the expressions are more involved as in the Gaussian case, but remain computationally tractable due to the restriction to isotropic jump amplitudes.
In the detailed implementation, I will consider $\mathbf{Y}(t)$ as a (generalized) Ornstein-Uhlenbeck process with $\mathbf{f}(\mathbf{x},t)=\frac{1}{2}\mathbf{x}$ and $g(t)=1$ such that Eq. (1) becomes
$\dot{\mathbf{Y}}(t)=-\frac{1}{2}\mathbf{Y}(t)+\boldsymbol{\xi}^{\rm G}(t)+\sum_{j=1}^{N_{T}}\mathbf{A}_{j}\delta(t-\tau_{j})$ (15)
In addition, I consider jumps with amplitudes $\mathbf{A}_{j}$ drawn from an isotropic multivariate Laplace distribution: $\mathbf{A}_{j}\sim\text{GL}(\sigma^{2},1)$, see Appendix D. This distribution not only represents a multivariate extension of the exponential distribution and is thus a standard case to consider beyond the Gaussian, but it also allows for straightforward sampling and has a simple characteristic function $G_{\rm L}(k)=(1+\frac{\sigma^{2}}{2}k^{2})^{-1}$, which simplifies the solution of the Ornstein-Uhlenbeck process Eq. (15). I refer to the resulting model as jump-diffusion Laplace (JDL) model. The propagator is then obtained by solving the Fokker-Planck equation associated with Eq. (15) leading to the solution (see Appendix C)
$p(\mathbf{x},t|\mathbf{x}^{\prime})$ $=$ $G_{1}(|\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}|,t)$ (16)
where $G_{1}(x,t)$ is given in terms of the inverse FT $G_{1}(x,t)=\mathcal{F}^{-1}\left\{\hat{G}_{1}(k,t)\right\}$ with
$\hat{G}_{1}(k,t)$ $=$ $\left(\frac{1+\frac{\sigma^{2}}{2}k^{2}e^{-t}}{1+\frac{\sigma^{2}}{2}k^{2}}\right)^{\lambda}e^{-\frac{D}{2}k^{2}\left(1-e^{-t}\right)}.$ (17)
Since $\hat{G}_{1}(k,t)$ is isotropic in $k$, evaluating $G_{1}(x,t)$ only requires a one-dimensional integral for arbitrary $d$. Likewise, the term $\nabla\mathcal{V}(\mathbf{x},t|\mathbf{x}^{\prime})$ in Eq. (12) can be evaluated by a one-dimensional integral due to isotropy
$\nabla\mathcal{V}(\mathbf{x},t|\mathbf{x}^{\prime})=\frac{\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}}{|\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}|}G_{2}(|\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}|,t)$ (18)
where
$G_{2}(x,t)=\frac{\partial}{\partial x}\mathcal{F}^{-1}\left\{\hat{G}_{1}(k,t)\frac{\psi(k)}{k^{2}}\right\}$ (19)
Explicit integral expressions for $G_{1,2}(x,t)$ are given in Eqs. (64,66). Evaluating these integrals is computationally costly, but since $\mathbf{f},g$ are typically unchanged for a given class of models (independent of $p_{\rm data}$), it is advantageous to create numerical representations of these functions offline and use them like analytical functions in different applications. This can be done by calculating the integrals at discretized points in the range $t\in[0,T]$ and $x\in[0,x_{\rm max}]$ and defining function approximations by interpolation.
From Eqs. (16,17) one obtains the FT of the stationary distribution $p_{\rm st}(x)=\lim_{t\to\infty}p(\mathbf{x},t|\mathbf{x}_{0})$ as
$\hat{p}_{\rm st}(k)=\left(1+\frac{\sigma^{2}}{2}k^{2}\right)^{-\lambda}e^{-\frac{D}{2}k^{2}}$ (20)
While the inverse FT is also here not available in closed form, it is straightforward to sample from $p_{\rm st}(x)$, see Eq. (25).
The JDL model implementation thus requires the following steps:
  1. 1.
    Evaluating the functions $G_{1,2}(x,t)$ numerically in terms of the integrals given in Eqs. (64,66) over the range $t\in[0,T]$ and $x\in[0,x_{\rm max}]$.
  2. 2.
    Optimizing Eq. (12) with the numerical representations of $p(\mathbf{x},t|\mathbf{x}^{\prime})$ and $\nabla\mathcal{V}(\mathbf{x},t|\mathbf{x}^{\prime})$ to obtain $\mathbf{s}_{\boldsymbol{\theta}}$ as approximation of the generalized score function. The Monte-Carlo integration requires simulation of Eq. (15), which is simplified due to the linear form of the dynamics. Direct integration yields
    $\mathbf{Y}(t)=\mathbf{Y}(0)e^{-t/2}+\bar{\boldsymbol{\xi}}_{-}^{\rm G}(t)+\bar{\mathbf{J}}_{-}(t),$ (21)
    where I define
    $\bar{\boldsymbol{\xi}}_{\pm}^{\rm G}(t)=\int_{0}^{t}\mathrm{d}s\,e^{\pm(t-s)/2}\boldsymbol{\xi}^{\rm G}(s),\qquad\qquad\bar{\mathbf{J}}_{\pm}(t)=\sum_{j=1}^{N_{t}}\mathbf{A}_{j}e^{\pm(t-\tau_{j})/2}.$ (22)
    Eq. (21) allows for exact sampling without discretization error, see Point 4 below for details.
  3. 3.
    Implementing the generative process via the probability flow ODE Eq. (10) or SDE Eq. (11) upon replacing the generalized score function by $\mathbf{s}_{\boldsymbol{\theta}}$. Efficient sampling can be implemented with exponential samplers [21] that exploit the linearity akin to Eq. (21), only that now a time discretization is needed due to the nonlinear drift. I use the simple discretizations:
    $\mathbf{X}_{i+1}=\mathbf{X}_{i}e^{\Delta t/2}+2\left(e^{\Delta t/2}-1\right)\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{X}_{i},T-t_{i})$ (23)
    for Eq. (10), and
    $\mathbf{X}_{i+1}=\mathbf{X}_{i}e^{\Delta t/2}+4\left(e^{\Delta t/2}-1\right)\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{X}_{i},T-t_{i})+\bar{\boldsymbol{\xi}}_{+}^{\rm G}(\Delta t)+\bar{\mathbf{J}}_{+}(\Delta t)$ (24)
    for Eq. (11). In Eqs. (23,24), the starting point is sampled $\mathbf{X}_{0}=\mathbf{X}(0)\sim p_{st}(x)$.
  4. 4.
    Sampling implementations:
    1. (a)
[…]
      where $\mathbf{Z}_{1}\sim\mathcal{N}(\mathbf{0},D\mathbf{I})$, $\mathbf{Z}_{2}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$, and $\Gamma\sim\text{Gamma}(\lambda)$.
    2. (b)
      Gaussian noise increments $\bar{\boldsymbol{\xi}}_{\pm}^{\rm G}(t)$. The definitions Eq. (22) imply $\bar{\boldsymbol{\xi}}_{-}^{\rm G}(t)\sim\mathcal{N}(\mathbf{0},D\left(1-e^{-t}\right))$ and $\bar{\boldsymbol{\xi}}_{+}^{\rm G}(t)\sim\mathcal{N}(\mathbf{0},D\left(e^{t}-1\right))$.
    3. (c)
[…]
4 Connection with LIM
It is instructive to compare the results in this work with those of the LIM framework of [20], where the forward process Eq. (1) is driven by $\alpha$-stable Lévy noise with $21<\alpha<$ 2. In this case, a modification of the standard score function has also been obtained leading to the fractional score function $\frac{\Delta^{\frac{\alpha-2}{2}}\nabla p(\mathbf{x},t)}{p(\mathbf{x},t)}$, where $\Delta^{\frac{\alpha-2}{2}}$ denotes the fractional Laplacian of order $\frac{\alpha-2}{2}$. While the derivation of the jump-diffusion generative process in Appendix A is not valid for this case, one can naively consider $\alpha$-stable noise in the generalized score function by substituting the corresponding $\psi(k)$ in Eq. (9). Indeed, the characteristic function of isotropic $\alpha$-stable noise increments $\Delta\boldsymbol{\xi}$ with $21<\alpha<$ 2 is likewise given by Eq. (7), but with $\psi(k)=k^{\alpha}$ [4]. For this special case, one finds from Eq. (9)
$\nabla\mathcal{V}(\mathbf{x},t)$ $=$ $g(t)^{\alpha}\nabla\mathcal{F}^{-1}\left\{\hat{p}(\mathbf{k},t)k^{\alpha-2}\right\}$ (26)
$=$ $g(t)^{\alpha}\Delta^{\frac{\alpha-2}{2}}\nabla p(\mathbf{x},t)$
and the generalized score function $\frac{\nabla\mathcal{V}(\mathbf{x},t)}{p(\mathbf{x},t)}$ becomes the fractional score function derived in [20]. The generative process Eq. (10) then recovers the probability flow formulation of LIM [20]. These results are somewhat surprising, since the rigorous analysis of the the time-reversal of Markov jump processes in [3] shows that in the case of $\alpha$-stable Lévy processes with $21<\alpha<$ 2, a very different time-reversal formula applies than for finite activity Lévy processes. However, the results for the generative process are consistent, which suggests that the generalized score function with $\mathcal{V}(\mathbf{x},t)$ of Eq. (9) might be valid more generally.
Specifying the forward process further as an Ornstein-Uhlenbeck process with $\mathbf{f}(\mathbf{x},t)=\frac{1}{2}\mathbf{x}$ and $g(t)=1$, it is also straightforward to show that the conditional generalized score function for $\psi(k)=k^{\alpha}$ has the form (see Appendix C.2)
$\frac{\nabla\mathcal{V}(\mathbf{x},t|\mathbf{x}^{\prime})}{p(\mathbf{x},t|\mathbf{x}^{\prime})}$ $=$ $-\frac{\left(\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}\right)}{2\left(1-e^{-t}\right)}$ (27)
recovering both the familiar Gaussian result for $\alpha=2$ and the result for the conditional fractional score function derived in [20] for $21<\alpha<$ 2 using a more cumbersome method. All power-law forms $\psi(k)\propto k^{\alpha}$ would thus lead to a simple functional form of the conditional generalized score function, if the validity of Eq. (9) could be extended to these cases.
5 Numerical experiment: heavy-tailed data distribution
(a)           (b) Method MSLE SE JDL-SDE 0.018 6.52×$\times 10^{-4}$ JDL-ODE 0.040 3.68×$\times 10^{-3}$ DLPM ($\alpha=1.7$) [14] 0.071 0.028 LIM ($\alpha=1.7$) [14] 0.267 0.077 DDPM [14] 0.528 0.4
Table 1: (a) Heavy-tailed samples from $p_{\rm data}$ given as an isotropic $\alpha$-stable distribution in $d=2$ as $\mathcal{F}^{-1}\left\{e^{-k^{\alpha}}\right\}$ with $\alpha=1.7$. Only one component is shown. (b) Measured values for the MSLE Eq. (28) over 20 runs. JDL-SDE and JDL-ODE values are calculated from 25,000 samples generated by Eqs. (23,24). The other values are taken from [14], whose $\alpha$-stable extension of DDPM is referred to as denoising Lévy probabilistic model (DLPM). For more details on the implementation, see Appendix F.
In order to demonstrate the applicability of the framework, I consider a simple toy experiment following [14], in which $p_{\rm data}$ is given by an $\alpha$-stable distribution with $\alpha=1.7$ in $d=2$. With such a strongly non-Gaussian data distribution one can test how well the generative process is able to capture heavy tails in the distribution, which is challenging with Gaussian methods. A suitable metric to compare the performance of different methods in approximating the tails is the mean-square logarithmic error (MSLE) of the higher quartiles, defined as [14]
${\rm MSLE}(\xi)=\int_{\xi}^{1}\mathrm{d}p\left(\log F_{\rm data}^{-1}(p)-\log F_{\rm gen}^{-1}(p)\right)^{2}$ (28)
where $\xi$ has been chosen as 0.95 and $F^{-1}$ denotes the inverse cumulative distribution function. For the implementation of JDL, Eq. (28) is evaluated separately for the two components and then averaged. As the results in Table 1 show, the JDL implementations in ODE and SDE form, Eqs. (23,24) strongly outperform both LIM [20] and DDPM with $\alpha$-stable noise [14]. In addition, the standard error of the samples is considerably smaller, by one magnitude for JDL-ODE and two magnitudes for JDL-SDE. It should be noted that neither the parameters of the jump-diffusion process (jump rate $\lambda$ and variance $\sigma^{2}$) nor the neural network used to estimate the generalized score function have been optimized in any way. Nevertheless, the JDL model provides better tail approximation than the two heavy-tailed methods with parameter $\alpha=1.7$, that is, set to the same stability index as the data distribution, where optimal performance might be expected. For more details on the implementation, see Appendix F
6 Conclusion
Extending the noise in the forward process to jump-diffusions leads to a straightforward non-Gaussian generalization of the work of [18]. The results highlight that the score function needs to be adapted to the noise and only takes the conventional form for the Gaussian noise case, while it is otherwise determined by the non-Gaussian details of the noise. Apart from investigating the performance of the framework in more realistic scenarios such as image synthesis, further detailed analysis is also needed to understand the effects of the various parameters and protocols underlying the implementation. One of the shortcomings of the approach is the representation of the generalized score function in terms of an inverse Fourier transform. Instead of using a numerical function approximation, it should also be possible to extend the sliced score matching approach [17] to the jump-diffusion case, which could provide an alternative approach to score estimation.
The fact that the JDL model outperforms LIM and DLPM on such a simple test despite not containing any heavy-tailed characteristics is intriguing. The somewhat poor performance of LIM might be related to the fact that the corresponding SDE neglects an additional noise term due to its finite variation [20]. As such noise and score-function are not exactly balanced as is required by the time reversal formalism and could lead to systematic discrepancies. Further investigation of this interesting theoretical aspect is planned in future work.
Appendix A Time-reversal of jump-diffusion processes
Before we look into more detail into the jump-diffusion generalization of the generative framework outlined in Sec. 2, a heuristic derivation of these results is presented. The aim is to derive an equation for the PDF $\tilde{p}(\mathbf{x},t)$ of the time-reverse or generative process. This PDF can be specified by the condition that $\tilde{p}(\mathbf{x},t)=p(\mathbf{x},T-t)$ for $t\in[0,T]$. Using the Fokker-Planck equation associated with Eq. (1) we obtain then
$\frac{\partial}{\partial t}\tilde{p}(\mathbf{x},t)$ $=$ $\frac{\partial}{\partial t}p(\mathbf{x},T-t)$ (29)
$=$ $-\left(-\nabla{\cdot}\mathbf{f}(\mathbf{X},T-t)p(\mathbf{x},T-t)+\frac{D}{2}g^{2}(T-t)\nabla^{2}p(\mathbf{x},T-t)\right)$
$=$ $\nabla{\cdot}\mathbf{f}(\mathbf{X},T-t)\tilde{p}(\mathbf{x},t)-\frac{D}{2}g^{2}(T-t)\nabla^{2}\tilde{p}(\mathbf{x},t)$
Eq. (29) can be expressed further in the form
$\frac{\partial}{\partial t}\tilde{p}(\mathbf{x},t)$ $=$ $\nabla{\cdot}\mathbf{f}(\mathbf{X},T-t)\tilde{p}(\mathbf{x},t)-Dg^{2}(T-t)\nabla^{2}\tilde{p}(\mathbf{x},t)+\frac{D}{2}g^{2}(T-t)\nabla^{2}\tilde{p}(\mathbf{x},t)$ (30)
$=$ $\nabla{\cdot}\left(\mathbf{f}(\mathbf{X},T-t)-Dg^{2}(T-t)\nabla\log\,\tilde{p}(\mathbf{x},t)\right)\tilde{p}(\mathbf{x},t)+\frac{D}{2}g^{2}(T-t)\nabla^{2}\tilde{p}(\mathbf{x},t)$
$=$ $\nabla{\cdot}\left(\mathbf{f}(\mathbf{x},T-t)-Dg^{2}(T-t)\nabla\log\,p(\mathbf{x},T-t)\right)\tilde{p}(\mathbf{x},t)$
$+\frac{D}{2}g^{2}(T-t)\nabla^{2}\tilde{p}(\mathbf{x},t)$
Eq. (30) is the Fokker-Planck equation corresponding to the SDE Eq. (2), which thus represents a stochastic process with PDF $\tilde{p}(\mathbf{x},t)$. The initial condition needs to satisfy $\mathbf{X}(0)\sim\tilde{p}(\mathbf{x},0)=p(\mathbf{x},T)$.
Rewriting the last term of Eq. (30) with $\tilde{p}(\mathbf{x},t)=p(\mathbf{x},T-t)$ yields further
$\frac{\partial}{\partial t}\tilde{p}(\mathbf{x},t)$ $=$ $\nabla{\cdot}\left(\mathbf{f}(\mathbf{x},T-t)-Dg^{2}(T-t)\nabla\log\,p(\mathbf{x},T-t)\right.$ (31)
$\left.+\frac{D}{2}g^{2}(T-t)\nabla\log\,p(\mathbf{x},T-t)\right)\tilde{p}(\mathbf{x},t)$
$=$ $-\nabla{\cdot}\left(-\mathbf{f}(\mathbf{x},T-t)+\frac{D}{2}g^{2}(T-t)\nabla\log\,p(\mathbf{x},T-t)\right)\tilde{p}(\mathbf{x},t)$
Eq. (31) is the Liouville equation corresponding to the deterministic dynamics Eq. (3), which likewise generates the PDF $\tilde{p}(\mathbf{x},t)$ for the initial condition $\mathbf{X}(0)\sim p(\mathbf{x},T)$. So we see that the generative process in the SDE description results from expressing the negative diffusion as a drift plus positive diffusion, while the ODE description results from expressing the negative diffusion as a drift only. We will see that the same heuristic prescription can also be applied for jump diffusions and agrees with rigorous results on the time-reversal of Markov processes with jumps, see Appendix E.
The forward process is now specified by Eq. (1) with noise $\boldsymbol{\xi}(t)$ characterized in Eqs. (6,7,8). The Fokker-Planck equation associated with $\mathbf{Y}(t)$ is obtained by standard methods [4] and has the integral form
$\frac{\partial}{\partial t}p(\mathbf{x},t)$ $=$ $-\nabla{\cdot}\mathbf{f}(\mathbf{x},t)p(\mathbf{x},t)+\frac{D}{2}g^{2}(t)\nabla^{2}p(\mathbf{x},t)$ (32)
$+\lambda\int\mathrm{d}\mathbf{z}\,\rho(z)\left[p(\mathbf{x}+g(t)\mathbf{z},t)-p(\mathbf{x},t)\right]$
The probability flow formalism for Lévy processes derived from Eq. (32) has been investigated in [8]. I likewise use the identity
$p(\mathbf{x}+g(t)\mathbf{z},t)-p(\mathbf{x},t)$ $=$ $\int_{0}^{1}\mathrm{d}\gamma\,\frac{\mathrm{d}}{\mathrm{d}\gamma}p(\mathbf{x}+\gamma g(t)\mathbf{z},t)$ (33)
$=$ $g(t)\int_{0}^{1}\mathrm{d}\gamma\,\mathbf{z}{\cdot}\nabla p(\mathbf{x}+\gamma g(t)\mathbf{z},t)$
such that Eq. (32) can be rewritten as
$\frac{\partial}{\partial t}p(\mathbf{x},t)$ $=$ $-\nabla{\cdot}\mathbf{f}(\mathbf{x},t)p(\mathbf{x},t)+\frac{D}{2}g^{2}(t)\nabla^{2}p(\mathbf{x},t)$ (34)
$+\lambda\,g(t)\int_{0}^{1}\mathrm{d}\gamma\int\mathrm{d}\mathbf{z}\,\rho(z)\mathbf{z}{\cdot}\nabla p(\mathbf{x}+\gamma g(t)\mathbf{z},t)$
$=$ $-\nabla{\cdot}\left(\mathbf{f}(\mathbf{x},t)-\frac{D}{2}g^{2}(t)\nabla\log\,p(\mathbf{x},t)\right.$
$\left.-\lambda\,g(t)\int_{0}^{1}\mathrm{d}\gamma\int\mathrm{d}\mathbf{z}\,\mathbf{z}\,\rho(z)\frac{p(\mathbf{x}+\gamma g(t)\mathbf{z},t)}{p(\mathbf{x},t)}\right)p(\mathbf{x},t)$
Starting again from the condition $\tilde{p}(\mathbf{x},t)=p(\mathbf{x},T-t)$, an evolution equation for $\tilde{p}$ is derived as
$\frac{\partial}{\partial t}\tilde{p}(\mathbf{x},t)$ $=$ $\frac{\partial}{\partial t}p(\mathbf{x},T-t)$ (35)
$=$ $\nabla{\cdot}\left(\mathbf{f}(\mathbf{x},T-t)-\frac{D}{2}g^{2}(T-t)\nabla\log\,p(\mathbf{x},T-t)\right.$
$\left.-\lambda\,g(T-t)\int_{0}^{1}\mathrm{d}\gamma\int\mathrm{d}\mathbf{z}\,\mathbf{z}\,\rho(z)\frac{p(\mathbf{x}+\gamma g(T-t)\mathbf{z},T-t)}{p(\mathbf{x},T-t)}\right)\tilde{p}(\mathbf{x},t)$
$=$ $-\nabla{\cdot}\left(-\mathbf{f}(\mathbf{x},T-t)+\frac{D}{2}g^{2}(T-t)\nabla\log\,p(\mathbf{x},T-t)\right.$
$\left.+\lambda\frac{\boldsymbol{\Lambda}(\mathbf{x},T-t)}{p(\mathbf{x},T-t)}\right)\tilde{p}(\mathbf{x},t)$
Here, I define the vector field
$\boldsymbol{\Lambda}(\mathbf{x},t)=g(t)\int_{0}^{1}\mathrm{d}\gamma\int\mathrm{d}\mathbf{z}\,\mathbf{z}\,\rho(z)p(\mathbf{x}+\gamma g(t)\mathbf{z},t)$ (36)
Below, I will derive an explicit expression for $\boldsymbol{\Lambda}$. Eq. (35) is the Liouville equation for the generative process, which takes the form of the probability flow equation
$\dot{\mathbf{X}}(t)=-\mathbf{f}(\mathbf{X},T-t)+\frac{D}{2}g^{2}(T-t)\nabla\log\,p(\mathbf{X},T-t)+\lambda\frac{\boldsymbol{\Lambda}(\mathbf{X},T-t)}{p(\mathbf{X},T-t)}$ (37)
The same result is obtained from rigorous results on the generator for time-reversed Markovian jump processes confirming the validity of Eq. (37), see Appendix E.
In order to keep a jump operator in the evolution equation for $\tilde{p}$ I simply introduce a factor 2 akin to the procedure in the Gaussian case. Eq. (35) can be accordingly rewritten as
$\frac{\partial}{\partial t}\tilde{p}(\mathbf{x},t)$ $=$ $-\nabla{\cdot}\left(-\mathbf{f}(\mathbf{x},T-t)+Dg^{2}(T-t)\nabla\log\,p(\mathbf{x},T-t)+2\lambda\frac{\boldsymbol{\Lambda}(\mathbf{x},T-t)}{p(\mathbf{x},T-t)}\right)\tilde{p}(\mathbf{x},t)$ (38)
$+\frac{D}{2}g^{2}(T-t)\nabla^{2}\log\,p(\mathbf{x},T-t)\tilde{p}(\mathbf{x},t)+\lambda\nabla\boldsymbol{\Lambda}(\mathbf{x},T-t)$
$=$ $-\nabla{\cdot}\left(-\mathbf{f}(\mathbf{x},T-t)+Dg^{2}(T-t)\nabla\log\,p(\mathbf{x},T-t)+2\lambda\frac{\boldsymbol{\Lambda}(\mathbf{x},T-t)}{p(\mathbf{x},T-t)}\right)\tilde{p}(\mathbf{x},t)$
$+\frac{D}{2}g^{2}(T-t)\nabla^{2}\tilde{p}(\mathbf{x},t)+\lambda\,g(T-t)\int_{0}^{1}\mathrm{d}\gamma\int\mathrm{d}\mathbf{z}\,\mathbf{z}\,\rho(z)\nabla\tilde{p}(\mathbf{x}+\gamma g(T-t)\mathbf{z},t)$
$=$ $-\nabla{\cdot}\left(-\mathbf{f}(\mathbf{x},T-t)+Dg^{2}(T-t)\nabla\log\,p(\mathbf{x},T-t)+2\lambda\frac{\boldsymbol{\Lambda}(\mathbf{x},T-t)}{p(\mathbf{x},T-t)}\right)\tilde{p}(\mathbf{x},t)$
$+\frac{D}{2}g^{2}(T-t)\nabla^{2}\tilde{p}(\mathbf{x},t)+\lambda\int\mathrm{d}\mathbf{z}\,\rho(z)\left[\tilde{p}(\mathbf{x}+g(T-t)\mathbf{z},t)-\tilde{p}(\mathbf{x},t)\right]$
using $\tilde{p}(\mathbf{x},t)=p(\mathbf{x},T-t)$ and Eq. (33) in the last step. Eq. (38) is the Fokker-Planck equation for the stochastic dynamics
$\dot{\mathbf{X}}(t)=-\mathbf{f}(\mathbf{X},T-t)+Dg^{2}(T-t)\nabla\log\,p(\mathbf{X},T-t)+2\lambda\frac{\boldsymbol{\Lambda}(\mathbf{X},T-t)}{p(\mathbf{X},T-t)}+g(T-t)\boldsymbol{\xi}(t)$ (39)
which likewise represents the generative process.
A.1 Generalized score function
In order to determine $\boldsymbol{\Lambda}$ further, I express $p$ as a Fourier transform and focus on a single component
$\Lambda_{i}(\mathbf{x},t)$ $=$ $g(t)\int_{0}^{1}\mathrm{d}\gamma\int\mathrm{d}\mathbf{z}\,z_{i}\,\rho(z)\,p(\mathbf{x}+\gamma g(t)\mathbf{z},t)$ (40)
$=$ $\frac{g(t)}{(2\pi)^{d}}\int\mathrm{d}\mathbf{k}\int_{0}^{1}\mathrm{d}\gamma\int\mathrm{d}\mathbf{z}\,z_{i}\,\rho(z)\,e^{-i\mathbf{k}{\cdot}(\mathbf{x}+\gamma g(t)\mathbf{z})}\hat{p}(\mathbf{k},t)$
$=$ $\frac{1}{(2\pi)^{d}}\int\mathrm{d}\mathbf{k}\,e^{-i\mathbf{k}{\cdot}\mathbf{x}}\hat{p}(\mathbf{k},t)\int_{0}^{1}\mathrm{d}\gamma\,\frac{i}{\gamma}\frac{\partial}{\partial k_{i}}\,\phi(k\,\gamma\,g(t)).$
In the last step I used the property
$\frac{\partial}{\partial k_{i}}\,\phi(k\,\gamma\,g(t))$ $=$ $\frac{\partial}{\partial k_{i}}\int\mathrm{d}\mathbf{z}\left(e^{-i\gamma g(t)\mathbf{k}{\cdot}\mathbf{z}}-1\right)\rho(z)$ (41)
$=$ $-i\gamma g(t)\int\mathrm{d}\mathbf{z}\,z_{i}\,\rho(z)e^{-i\gamma g(t)\mathbf{k}{\cdot}\mathbf{z}}$
due to Eq. (8) and isotropy. This yields further, reminding that $k=\|\mathbf{k}\|=\left(\sum_{i=1}^{d}k_{i}^{2}\right)^{1/2}$
$\int_{0}^{1}\mathrm{d}\gamma\,\frac{i}{\gamma}\frac{\partial}{\partial k_{i}}\,\phi(k\gamma g(t))$ $=$ $\int_{0}^{1}\mathrm{d}\gamma\,ig(t)\,\phi^{\prime}(k\,\gamma\,g(t))\frac{\partial k}{\partial k_{i}}$ (42)
$=$ $i\,\phi(kg(t))\frac{1}{k}\frac{\partial k}{\partial k_{i}}$
$=$ $\frac{i\,k_{i}}{k^{2}}\phi(kg(t))$
Substituting back into Eq. (40) yields the result for $\Lambda_{i}$ as an inverse Fourier transform
$\Lambda_{i}(\mathbf{x},t)$ $=$ $-\frac{\partial}{\partial x_{i}}\frac{1}{(2\pi)^{d}}\int\mathrm{d}\mathbf{k}\,e^{-i\mathbf{k}{\cdot}\mathbf{x}}\hat{p}(\mathbf{k},t)\frac{\phi(k\,g(t))}{k^{2}}.$ (43)
A notational simplification can be introduced by defining (cf. Eq. (9))
$\mathcal{V}(\mathbf{x},t)$ $=$ $\frac{D}{2}g^{2}(t)\,p(\mathbf{x},t)-\frac{\lambda}{(2\pi)^{d}}\int\mathrm{d}\mathbf{k}\,e^{-i\mathbf{k}{\cdot}\mathbf{x}}\hat{p}(\mathbf{k},t)\frac{\phi(k\,g(t))}{k^{2}}$ (44)
$=$ $\frac{1}{(2\pi)^{d}}\int\mathrm{d}\mathbf{k}\,e^{-i\mathbf{k}{\cdot}\mathbf{x}}\hat{p}(\mathbf{k},t)\frac{\psi(k\,g(t))}{k^{2}}$
where $\psi(k)$ is defined in Eq. (7). With Eq. (43), we can write the generative process Eq. (37) as
$\dot{\mathbf{X}}(t)=-\mathbf{f}(\mathbf{X},T-t)+\frac{\nabla\mathcal{V}(\mathbf{X},T-t)}{p(\mathbf{X},T-t)}$ (45)
and likewise in Eq. (39), leading to Eqs. (10,11). The generalized score function is thus identified as $\frac{\nabla\mathcal{V}(\mathbf{x},t)}{p(\mathbf{x},t)}$, which is the central quantity in the jump-diffusion approach to generative modelling.
Appendix B Denoising score matching loss function
In order to derive the jump-diffusion equivalent of the denoising score matching (DSM) loss function Eq. (4), I start with the explicit score matching (ESM) loss function for a given weight function $w(t)$ and follow the derivation of the conventional DSM loss function in [19]:
$\mathcal{J}_{\rm ESM}(\boldsymbol{\theta})$ $=$ $\int_{0}^{T}\mathrm{d}t\,w(t)\,\mathbb{E}_{\mathbf{Y}(t)}\left[\left\|\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{Y},t)-\frac{\nabla\mathcal{V}(\mathbf{Y},t)}{p(\mathbf{Y},t)}\right\|^{2}\right]$ (46)
$=$ $\int_{0}^{T}\mathrm{d}t\,w(t)\,\mathbb{E}_{\mathbf{Y}(t)}\left[\|\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{Y},t)\|^{2}-2\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{Y},t){\cdot}\frac{\nabla\mathcal{V}(\mathbf{Y},t)}{p(\mathbf{Y},t)}+\left\|\frac{\nabla\mathcal{V}(\mathbf{Y},t)}{p(\mathbf{Y},t)}\right\|^{2}\right]$
$=$ $\int_{0}^{T}\mathrm{d}t\,w(t)\left(\mathbb{E}_{\mathbf{Y}(t)}\left[\|\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{Y},t)\|^{2}\right]-2\,\mathbb{E}_{\mathbf{Y}(t)}\left[\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{Y},t){\cdot}\frac{\nabla\mathcal{V}(\mathbf{Y},t)}{p(\mathbf{Y},t)}\right]\right)$
$+C_{1}(T)$
since the last term is independent of $\boldsymbol{\theta}$. For the second term I obtain further
$\mathbb{E}_{\mathbf{Y}(t)}\left[\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{Y},t){\cdot}\frac{\nabla\mathcal{V}(\mathbf{Y},t)}{p(\mathbf{Y},t)}\right]$ $=$ $\int\mathrm{d}\mathbf{x}\,\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{x},t){\cdot}\frac{\nabla\mathcal{V}(\mathbf{x},t)}{p(\mathbf{x},t)}p(\mathbf{x},t)$ (47)
$=$ $\int\mathrm{d}\mathbf{x}\,\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{x},t){\cdot}\nabla\mathcal{V}(\mathbf{x},t)$
$=$ $\int\mathrm{d}\mathbf{x}\,\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{x},t){\cdot}\nabla\int\mathrm{d}\mathbf{x}^{\prime}\mathcal{V}(\mathbf{x},t|\mathbf{x}^{\prime})p_{\rm data}(\mathbf{x}^{\prime})$
In analogy to Eq. (44) I define now (cf. Eq. (13))
$\mathcal{V}(\mathbf{x},t|\mathbf{x}^{\prime})$ $=$ $\frac{1}{(2\pi)^{d}}\int\mathrm{d}\mathbf{k}\,e^{-i\mathbf{k}{\cdot}\mathbf{x}}\hat{p}(\mathbf{k},t|\mathbf{x}^{\prime})\frac{\psi(k\,g(t))}{k^{2}}$ (48)
where $\hat{p}(\mathbf{k},t|\mathbf{x}^{\prime})$ denotes the FT of the propagator $p(\mathbf{x},t|\mathbf{x}^{\prime})$. Continuing from Eq. (47) yields
$\mathbb{E}_{\mathbf{Y}(t)}\left[\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{Y},t){\cdot}\frac{\nabla\mathcal{V}(\mathbf{Y},t)}{p(\mathbf{Y},t)}\right]$ $=$ $\int\mathrm{d}\mathbf{x}\int\mathrm{d}\mathbf{x}^{\prime}\,\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{x},t){\cdot}\nabla\mathcal{V}(\mathbf{x},t|\mathbf{x}^{\prime})\frac{p(\mathbf{x},t|\mathbf{x}^{\prime})}{p(\mathbf{x},t|\mathbf{x}^{\prime})}p_{\rm data}(\mathbf{x}^{\prime})$ (49)
$=$ $\int\mathrm{d}\mathbf{x}\int\mathrm{d}\mathbf{x}^{\prime}\,\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{x},t){\cdot}\frac{\nabla\mathcal{V}(\mathbf{x},t|\mathbf{x}^{\prime})}{p(\mathbf{x},t|\mathbf{x}^{\prime})}p(\mathbf{x},t|\mathbf{x}^{\prime})p_{\rm data}(\mathbf{x}^{\prime})$
$=$ $\mathbb{E}_{\mathbf{Y}(0)}\mathbb{E}_{\mathbf{Y}(t)|\mathbf{Y}(0)}\left[\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{Y},t){\cdot}\frac{\nabla\mathcal{V}(\mathbf{Y},t|\mathbf{Y}_{0})}{p(\mathbf{Y},t|\mathbf{Y}_{0})}\right]$
The equivalent of Eq. (4) is thus $\mathcal{J}_{\rm JD}$ of Eq. (12), which satisfies
$\mathcal{J}_{\rm JD}(\boldsymbol{\theta})$ $=$ $\int_{0}^{T}\mathrm{d}t\,w(t)\left(\mathbb{E}_{\mathbf{Y}(t)}\left[\|\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{Y},t)\|^{2}\right]\right.$ (50)
$\left.-2\,\mathbb{E}_{\mathbf{Y}(0)}\mathbb{E}_{\mathbf{Y}(t)|\mathbf{Y}(0)}\left[\mathbf{s}_{\boldsymbol{\theta}}(\mathbf{Y},t){\cdot}\frac{\nabla\mathcal{F}(\mathbf{Y},t|\mathbf{Y}_{0})}{p(\mathbf{Y},t|\mathbf{Y}_{0})}\right]\right)+C_{2}(T)$
$=$ $J_{\rm ESM}(\boldsymbol{\theta})-C_{1}(T)+C_{2}(T)$
due to Eq. (49) and thus optimizing $\mathcal{J}_{\rm JD}(\boldsymbol{\theta})$ with respect to $\boldsymbol{\theta}$ is equivalent to optimizing $\mathcal{J}_{\rm ESM}(\boldsymbol{\theta})$.
Appendix C The generalized Ornstein-Uhlenbeck process
The forward process is given by the generalized Ornstein-Uhlenbeck process Eq. (15). I first consider an arbitrary jump amplitude distribution $\rho(z)$ with corresponding $\phi(k)$, Eq. (8). The Fokker-Planck equation associated with Eq. (15) is given by setting $\mathbf{f}(\mathbf{x},t)=\frac{1}{2}$ and $g(t)=1$ in Eq. (32), which we want to solve for the conditional PDF $p(\mathbf{x},t|\mathbf{x}^{\prime})$ with the initial condition $p(\mathbf{x},0|\mathbf{x}^{\prime})=\delta(\mathbf{x}-\mathbf{x}^{\prime})$. Fourier-transforming this Fokker-Planck equation yields
$\frac{\partial}{\partial t}\hat{p}(\mathbf{k},t|\mathbf{x}^{\prime})$ $=$ $\left[-\frac{1}{2}\mathbf{k}{\cdot}\nabla_{\mathbf{k}}-\frac{D}{2}k^{2}+\lambda\phi(k)\right]\hat{p}(\mathbf{k},t|\mathbf{x}^{\prime})$ (51)
$=$ $\left[-\frac{1}{2}\mathbf{k}{\cdot}\nabla_{\mathbf{k}}-\psi(k)\right]\hat{p}(\mathbf{k},t|\mathbf{x}^{\prime})$
with initial condition $\hat{p}(\mathbf{k},0|\mathbf{x}^{\prime})=e^{i\mathbf{k}{\cdot}\mathbf{x}^{\prime}}$. This first-order equation can be solved with the method of characteristics, which yields the solution
$\hat{p}(\mathbf{k},t|\mathbf{x}^{\prime})$ $=$ $\exp\left\{i\mathbf{k}{\cdot}\mathbf{x}^{\prime}e^{-t/2}-\int_{0}^{t}\mathrm{d}u\,\psi(k\,e^{-u/2})\right\}$ (52)
$=$ $e^{i\mathbf{k}{\cdot}\mathbf{x}^{\prime}e^{-t/2}}\hat{G}_{1}(k,t)$
where
$\hat{G}_{1}(k,t)$ $=$ $\exp\left\{-\int_{0}^{t}\mathrm{d}u\,\psi(k\,e^{-u/2})\right\}$ (53)
$=$ $\exp\left\{-\frac{D}{2}k^{2}\left(1-e^{-t}\right)+\lambda\int_{0}^{t}\mathrm{d}u\,\phi(k\,e^{-u/2})\right\}.$ (54)
The propagator is given as the inverse FT of $\hat{p}(\mathbf{k},0|\mathbf{x}^{\prime})$ which can be written in the form
$p(\mathbf{x},t|\mathbf{x}^{\prime})$ $=$ $G_{1}(|\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}|,t)$ (55)
with the isotropic function
$G_{1}(x,t)$ $=$ $\frac{1}{(2\pi)^{d}}\int\mathrm{d}\mathbf{k}\,e^{-i\mathbf{k}{\cdot}\mathbf{x}}\hat{G}_{1}(k,t).$ (56)
An expression for $\mathcal{V}(\mathbf{x},t|\mathbf{x}^{\prime})$ can then be directly derived from the definition Eq. (13), for which the FT reads simply
$\hat{\mathcal{V}}(\mathbf{k},t|\mathbf{x}^{\prime})$ $=$ $\hat{p}(\mathbf{k},t|\mathbf{x}^{\prime})\frac{\psi(k)}{k^{2}}$ (57)
$=$ $e^{i\mathbf{k}{\cdot}\mathbf{x}^{\prime}e^{-t/2}}\hat{G}_{1}(k,t)\frac{\psi(k)}{k^{2}}$
due to Eq. (52). Taking the gradient of the inverse FT yields
$\nabla\mathcal{V}(\mathbf{x},t|\mathbf{x}^{\prime})$ $=$ $\frac{1}{(2\pi)^{d}}\nabla\int\mathrm{d}\mathbf{k}\,e^{-i\mathbf{k}{\cdot}\left(\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}\right)}\hat{G}_{1}(k,t)\frac{\psi(k)}{k^{2}}$ (58)
$=$ $\frac{\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}}{|\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}|}G_{2}(|\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}|,t)$ (59)
upon defining the isotropic function
$G_{2}(x,t)=\frac{1}{(2\pi)^{d}}\frac{\partial}{\partial x}\int\mathrm{d}\mathbf{k}\,e^{-i\mathbf{k}{\cdot}\mathbf{x}}\hat{G}_{1}(k,t)\frac{\psi(k)}{k^{2}}$ (60)
C.1 Laplace distributed amplitudes
In order to simplify the framework, it is desirable to have a closed analytic form for $\hat{p}(\mathbf{k},0|\mathbf{x}^{\prime})$, which is achieved, for example, by choosing amplitudes according to an isotropic multivariate Laplace distribution $\rho=\text{GL}(0,\sigma^{2}\mathbf{I},1)$, see Appendix D for further details on the multivariate Laplace distribution. The function $\phi(k)$ of Eq. (8) becomes then
$\phi_{\rm L}(k)=\frac{1}{1+\frac{\sigma^{2}}{2}k^{2}}-1$ (61)
and
$\int_{0}^{t}\mathrm{d}u\,\phi_{\rm L}(k\,e^{-u/2})=\log\left(\frac{1+\frac{\sigma^{2}}{2}k^{2}e^{-s}}{1+\frac{\sigma^{2}}{2}k^{2}}\right)$ (62)
Eq. (53) becomes
$\hat{G}_{1}(k,t)=\left(\frac{1+\frac{\sigma^{2}}{2}k^{2}e^{-t}}{1+\frac{\sigma^{2}}{2}k^{2}}\right)^{\lambda}e^{-\frac{D}{2}k^{2}\left(1-e^{-t}\right)}$ (63)
After integrating out the orientational degrees of freedom, the function $G_{1}(x,t)$, Eq. (56), can be expressed as the one-dimensional integral
$G_{1}(x,t)$ $=$ $\frac{1}{(2\pi)^{d/2}}\int_{0}^{\infty}\mathrm{d}k\,k\,\left(\frac{k}{x}\right)^{\frac{d}{2}-1}J_{\frac{d}{2}-1}(k\,x)\left(\frac{1+\frac{\sigma^{2}}{2}k^{2}e^{-t}}{1+\frac{\sigma^{2}}{2}k^{2}}\right)^{\lambda}e^{-\frac{D}{2}k^{2}\left(1-e^{-t}\right)}$ (64)
where $J_{\mu}$ denotes the Bessel function of the first kind. For $t\to\infty$ the stationary distribution $p_{\rm st}(x)=\lim_{t\to\infty}p(\mathbf{x},t|\mathbf{x}_{0})$ is obtained as
$p_{\rm st}(x)$ $=$ $\lim_{t\to\infty}G_{1}(x,t)$ (65)
$=$ $\frac{1}{(2\pi)^{d}}\int\mathrm{d}\mathbf{k}\,e^{-i\mathbf{k}{\cdot}\mathbf{x}}\left(\frac{1}{1+\frac{\sigma^{2}}{2}k^{2}}\right)^{\lambda}e^{-\frac{D}{2}k^{2}}$
The function $G_{2}(x,t)$, Eq. (60), can likewise be expressed as a one-dimensional integral:
$G_{2}(x,t)$ $=$ $\frac{1}{(2\pi)^{d/2}}\frac{\partial}{\partial x}\int_{0}^{\infty}\mathrm{d}k\,k\,\left(\frac{k}{x}\right)^{\frac{d}{2}-1}J_{\frac{d}{2}-1}(k\,x)\hat{G}_{1}(k,t)\frac{\psi(k)}{k^{2}}$ (66)
$=$ $-\frac{x}{(2\pi)^{d/2}}\int_{0}^{\infty}\mathrm{d}k\,k\,\left(\frac{k}{x}\right)^{\frac{d}{2}}J_{\frac{d}{2}}(k\,x)\left(\frac{1+\frac{\sigma^{2}}{2}k^{2}e^{-t}}{1+\frac{\sigma^{2}}{2}k^{2}}\right)^{\lambda}\times$
$e^{-\frac{D}{2}k^{2}\left(1-e^{-t}\right)}\left(\frac{D}{2}-\lambda\frac{1}{k^{2}\left(1+\frac{\sigma^{2}}{2}k^{2}\right)}+1\right)$
substituting $\psi(k)$ of Eq. (7) with Eq. (61).
C.2 The conditional generalized score function for $\psi(k)=k^{\alpha}$
With Eqs. (58,53) one obtains for the special case $\psi(k)=k^{\alpha}$
$\nabla\mathcal{V}(\mathbf{x},t|\mathbf{x}^{\prime})$ $=$ $\frac{1}{(2\pi)^{d}}\nabla\int\mathrm{d}\mathbf{k}\,e^{-i\mathbf{k}{\cdot}\left(\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}\right)}e^{-k^{\alpha}f(t)}k^{\alpha-2}$ (67)
$=$ $\frac{1}{(2\pi)^{d}}\int\mathrm{d}\mathbf{k}\,e^{-i\mathbf{k}{\cdot}\left(\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}\right)}(-i\mathbf{k})e^{-k^{\alpha}f(t)}k^{\alpha-2}$
$=$ $\frac{1}{(2\pi)^{d}}\int\mathrm{d}\mathbf{k}\,e^{-i\mathbf{k}{\cdot}\left(\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}\right)}i\nabla_{\mathbf{k}}\frac{e^{-k^{\alpha}f(t)}}{\alpha\,f(t)}$
$=$ $-\frac{\left(\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}\right)}{\alpha\,f(t)}\frac{1}{(2\pi)^{d}}\int\mathrm{d}\mathbf{k}\,e^{-i\mathbf{k}{\cdot}\left(\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}\right)}e^{-k^{\alpha}f(t)}$
$=$ $-\frac{\left(\mathbf{x}-\mathbf{x}^{\prime}e^{-t/2}\right)}{\alpha\,f(t)}p(\mathbf{x},t|\mathbf{x}^{\prime})$
where $f(t)=\int_{0}^{t}\mathrm{d}u\,e^{-\alpha\,u/2}=\frac{2}{\alpha}\left(1-e^{-t}\right)$. Thus the conditional generalized score function becomes Eq. (27).
Appendix D The multivariate generalized Laplace distribution
I briefly review basic properties of the multivariate generalized Laplace distribution following [9].
Definition (Multivariate generalized Laplace law). A random vector in $\mathbb{R}^{d}$ is said to have a multivariate generalized asymmetric Laplace distribution, if its characteristic function is given by
$G(\mathbf{k})=\left(\frac{1}{1+\frac{1}{2}\mathbf{k}^{\rm T}\boldsymbol{\Sigma}\mathbf{k}-i\boldsymbol{\mu}{\cdot}\mathbf{k}}\right)^{\nu},\qquad\qquad\mathbf{k}\in\mathbb{R}^{d}$ (68)
where $\nu>0$, $\boldsymbol{\mu}\in\mathbb{R}^{d}$, and $\boldsymbol{\Sigma}$ is a $d\times d$ non-negative definite symmetric matrix. This distribution is denoted by $\text{GAL}(\boldsymbol{\mu},\boldsymbol{\Sigma},\nu)$.
Two relevant properties:
  1. 1.
    A random vector $\mathbf{L}\sim\text{GAL}(\boldsymbol{\mu},\boldsymbol{\Sigma},\nu)$ can be interpreted as a subordinated Gaussian process such that
[…]
    where $\mathbf{X}\sim\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma})$ and $\Gamma$ has a standard Gamma distribution with shape parameter $\nu$.
  2. 2.
    The mean vector and covariance matrix of $\mathbf{L}\sim\text{GAL}(\boldsymbol{\mu},\boldsymbol{\Sigma},\nu)$ are given by
    $\mathbb{E}(\mathbf{L})=\boldsymbol{\mu}\,\nu,\qquad\qquad\text{Cov}(\mathbf{L})=\nu(\boldsymbol{\Sigma}+\boldsymbol{\mu}{\cdot}\boldsymbol{\mu})$ (70)
Since we are interested in describing the jump amplitudes in the noise process $\boldsymbol{\xi}(t)$, Eq. (6), we want no contribution to the drift leading to the constraint $\boldsymbol{\mu}=\mathbf{0}$. We also require an isotropic distribution, which is realized by setting $\boldsymbol{\Sigma}=\sigma^{2}\mathbf{I}$. The resulting distribution can be called the isotropic multivariate generalized Laplace distribution $\text{GL}(\sigma^{2},\nu)=\text{GAL}(\mathbf{0},\sigma^{2}\mathbf{I},\nu)$ and has the characteristic function
$G(k)=\left(1+\frac{\sigma^{2}}{2}k^{2}\right)^{-\nu}$ (71)
Amplitude samples can be easily generated using Eq. (69): $\mathbf{L}\sim\text{GL}(\sigma^{2},\nu)$ has the representation
[…]
where $\mathbf{X}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$ and $\Gamma\sim\text{Gamma}(\nu)$.
Appendix E Alternative derivation of the Liouville equation for $\tilde{p}$
The time-reversal of Markovian jump processes has been rigorously investigated in [3]. Considering the case $D=0$, the generator of the time-reversal of $\mathbf{Y}(t)$, Eqs. (1,7,8), is shown to have the form
~$\tilde{\mathcal{L}}u(\mathbf{x})$ $=$ $-\mathbf{f}(\mathbf{x},T-t){\cdot}\nabla u(\mathbf{x})$ (73)
$+\lambda\int\mathrm{d}\mathbf{z}\,\rho(z)\left[u(\mathbf{x}+g(T-t)\mathbf{z})-u(\mathbf{x})\right]\frac{p(\mathbf{x}+g(T-t)\mathbf{z},T-t)}{p(\mathbf{x},T-t)}$
where $p(\mathbf{x},t)$ is the PDF of $\mathbf{Y}(t)$. Using the same identity as in Eq. (33) leads to
~$\tilde{\mathcal{L}}u(\mathbf{x})$ $=$ $-\mathbf{f}(\mathbf{x},T-t){\cdot}\nabla u(\mathbf{x})+\lambda\,g(T-t)\int\mathrm{d}\mathbf{z}\,\rho(z)\int_{0}^{1}\mathrm{d}\gamma\,\left[\mathbf{z}{\cdot}\nabla u(\mathbf{x}+\gamma g(T-t)\mathbf{z})\right]\times$ (74)
$\frac{p(\mathbf{x}+g(T-t)\mathbf{z},T-t)}{p(\mathbf{x},T-t)}$
This equation can be used to determine the adjoint generator $\tilde{\mathcal{L}}^{\dagger}$ that governs the time evolution of $\tilde{p}(\mathbf{x},t)$, the PDF of the time-reversed process $\mathbf{X}(t)$:
$\int\mathrm{d}\mathbf{x}\,\tilde{p}(\mathbf{x},t)\tilde{\mathcal{L}}u(\mathbf{x})$ $=$ $-\int\mathrm{d}\mathbf{x}\,\tilde{p}(\mathbf{x},t)\mathbf{f}(\mathbf{x},T-t){\cdot}\nabla u(\mathbf{x})+\lambda\,g(T-t)\int\mathrm{d}\mathbf{x}\,\tilde{p}(\mathbf{x},t)\times$ (75)
$\int_{0}^{1}\mathrm{d}\gamma\int\mathrm{d}\mathbf{z}\,\rho(z)\,\left[\mathbf{z}{\cdot}\nabla u(\mathbf{x}+\gamma g(T-t)\mathbf{z})\right]\frac{p(\mathbf{x}+g(T-t)\mathbf{z},T-t)}{p(\mathbf{x},T-t)}$
$=$ $\int\mathrm{d}\mathbf{x}\,u(\mathbf{x})\nabla{\cdot}\mathbf{f}(\mathbf{x},T-t)\tilde{p}(\mathbf{x},t)+\lambda\,g(T-t)\int\mathrm{d}\mathbf{x}\int_{0}^{1}\mathrm{d}\gamma\int\mathrm{d}\mathbf{z}\times$
$\left[\mathbf{z}{\cdot}\nabla u(\mathbf{x}+\gamma g(T-t)\mathbf{z})\right]p(\mathbf{x}+g(T-t)\mathbf{z},T-t)$
$=$ $\int\mathrm{d}\mathbf{x}\,u(\mathbf{x})\nabla{\cdot}\mathbf{f}(\mathbf{x},T-t)\tilde{p}(\mathbf{x},t)-\lambda\,g(T-t)\int\mathrm{d}\mathbf{x}\,u(\mathbf{x})\times$
$\int_{0}^{1}\mathrm{d}\gamma\int\mathrm{d}\mathbf{z}\,\rho(z)\,\mathbf{z}{\cdot}\nabla p(\mathbf{x}+(1-\gamma)g(T-t)\mathbf{z},T-t)$
$=$ $\int\mathrm{d}\mathbf{x}\,u(\mathbf{x})\nabla{\cdot}\mathbf{f}(\mathbf{x},T-t)\tilde{p}(\mathbf{x},t)-\lambda\,g(T-t)\int\mathrm{d}\mathbf{x}\,u(\mathbf{x})\times$
$\int_{0}^{1}\mathrm{d}\gamma\int\mathrm{d}\mathbf{z}\,\rho(z)\,\mathbf{z}{\cdot}\nabla\,\frac{p(\mathbf{x}+\gamma g(T-t)\mathbf{z},T-t)}{p(\mathbf{x},T-t)}\tilde{p}(\mathbf{x},t)$
using $\tilde{p}(\mathbf{x},t)=p(\mathbf{x},T-t)$. Thus the adjoint generator is given by
~$\tilde{\mathcal{L}}^{\dagger}\tilde{p}(\mathbf{x},t)$ $=$ $-\nabla{\cdot}\left(-\mathbf{f}(\mathbf{x},T-t)+\lambda\,g(T-t)\int_{0}^{1}\mathrm{d}\gamma\int\mathrm{d}\mathbf{z}\,\mathbf{z}\,\rho(z)\right.\times$ (76)
$\left.\frac{p(\mathbf{x}+\gamma g(T-t)\mathbf{z},T-t)}{p(\mathbf{x},T-t)}\right)\tilde{p}(\mathbf{x},t).$
In the presence of the additional Gaussian noise component in Eq. (7) for $D\neq 0$ one simply adds the standard diffusive term of Eq. (29). This yields Eq. (35) for the time-evolution of $\tilde{p}$.
Appendix F Numerical experiment
The implementation of the test is done in analogy to that in [14] so that values are comparable. The only difference is that the MSLE values for JDL-SDE and JDL-ODE are averaged over the two components, while the values from [14] are evaluated over the first component only. Due to the isotropy of $p_{\rm data}$ the two MSLE procedures lead to comparable values. 32,000 samples from $p_{\rm data}$ have been used for estimation of the generalized score function and 25,000 samples have been generated for testing. The functions $G_{1,2}(x,t)$ that determine the generalized score function via Eqs. (16,18) are evaluated by discretizing $t\in[0,10]$ and $x\in[0,10]$ into 200 equally spaced values and creating a function representation by interpolation. The neural network parametrizing $\mathbf{s}_{\boldsymbol{\theta}}$ has been kept simple, consisting of 4 layers with 200 elementwise $\tanh$ activation nodes. The weight in the loss Eq. (12) is taken as $w(t)=1/T$. The JDL implementation in Eqs. (21,23,24) uses $D=1$, $\lambda=1$, $\sigma^{2}=2$, $T=10$, $\Delta t=0.1$. For further details on the implementations of DLPM, LIM, DDPM, please see [14].